Suppose x has a distribution with μ = 10 and σ = 7.

(a) If a random sample of size n = 40 is drawn, find μx, σ x and P(10 ≤ x ≤ 12). (Round σx to two decimal places and the probability to four decimal places.)

μx =
σx =
P(10 ≤ x ≤ 12) =


(b) If a random sample of size n = 63 is drawn, find μx, σ x and P(10 ≤ x ≤ 12). (Round σ x to two decimal places and the probability to four decimal places.)

μx =
σx =
P(10 ≤ x ≤ 12) =


(c) Why should you expect the probability of part (b) to be higher than that of part (a)? (Hint: Consider the standard deviations in parts (a) and (b).)
The standard deviation of part (b) is
larger than / the same as / smaller than part (a) because of the smaller / same / larger sample size. Therefore, the distribution about μx is wider / the same / narrower.
Answer #1

solution Gives el = 10 6 = 7 a) D:40 e x = 10 P (losxs12) = PC 10-10 < x- el x < 12-10 .11 6x = P( 525 22 ) =P(0323 1.80) = Pb) n = 63 el x = 10 6xCo - More - 0.88 p(105x512) =P(10-10 X-el x x 12-101 0.88 6x 0.88 2. 0.88525 ose) = poszs 2.27) = P(ZC